OpenStudy (anonymous):

Find all points having an x-coordinate of 7 whose distance from the point (4,-2) is 5

OpenStudy (mertsj):

Write the equation of a circle with center (4,-2) and radius 5. Then replace x with 7 and solve for y.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let the point be (7,y) \[\sqrt{\left( 7-4 \right)^{2}+\left( y+2 \right)^{2}}=5,\left( y+2 \right)^{2}=25-9=16\] \[y +2=\pm 4\] solve for y and get the two values of y. Then the points are (7,y) and (7,-y)

OpenStudy (anonymous):

correction y means one value -y means second value